OpenStudy (anonymous):

Could someone show me how to do this and the formula? Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3

OpenStudy (nikato):

S the standard form of a circle is (x-h)^2 + (x-k)^2 = r^2

OpenStudy (nikato):

Where h,k is the center and r= radius

OpenStudy (anonymous):

(x-5)^2 + (y-6)^2 = 3^2

OpenStudy (anonymous):

Okay, so, @adrdiazes is that my question plugged into the formula?

OpenStudy (nikato):

Yea, so that would the equation for ur circle

OpenStudy (anonymous):

So is that my answer or do I do anything else to it? :x

OpenStudy (nikato):

That's should be it

OpenStudy (anonymous):

This is your answer ---> (x-5)^2 + (y-6)^2 = 3^2 :D

OpenStudy (anonymous):

Oh okay, that was easy! Thank you both so much! <3

OpenStudy (anonymous):

You're welcome :)

OpenStudy (nikato):

Your welcome

OpenStudy (anonymous):

So this is correct, right? @nikato @AdrDiazEs

OpenStudy (anonymous):

YEs, that's good :)

OpenStudy (anonymous):

Phew! Okay, thank you!